OpenStudy (anonymous):

e√5 Exact value of expression? Logorithms

OpenStudy (anonymous):

Answers:(A) √5 (B) 5 (C)e (D)25

OpenStudy (anonymous):

**Ln e^√5

jimthompson5910 (jim_thompson5910):

\[\Large \ln\left(e^{\sqrt{5}}\right)\] \[\Large \sqrt{5}\ln\left(e\right)\] \[\Large \sqrt{5}(1)\] \[\Large \sqrt{5}\] So \[\Large \ln\left(e^{\sqrt{5}}\right)=\sqrt{5}\]

OpenStudy (slaaibak):

A

OpenStudy (anonymous):

Thanks!